OpenStudy (anonymous):

Find the domain on the function. Write your answer in interval notation. If the answer includes more than one interval write the intervals separated by the union symbol, U.

OpenStudy (anonymous):

f(x)=sqrt(73.5-7x)

OpenStudy (noelgreco):

What limits the domain when you have an even root? In other words, what has to be true of the radicand, assuming the domain is the set of real numbers?

OpenStudy (anonymous):

I have no idea how to even start... Idk...

OpenStudy (anonymous):

It cannot be zero right?

OpenStudy (noelgreco):

Bingo! How can you figure out the value of x that makes the radicand (the expression under the radical) = 0?

OpenStudy (noelgreco):

Actually, it can be zero, it can't be negative.

OpenStudy (noelgreco):

73.5 - 7x = 0. Solve, but you're not done.

hero (hero):

\(\sqrt{0} = 0\)

OpenStudy (anonymous):

do I solve 73.5-7x=0?

hero (hero):

You find the domain by solving \(73.5 - 7x \ge 0\)

OpenStudy (noelgreco):

It's easiest to solve the equality and then consider whether an increasing or decreasing value of x satisfies the inequality.